The numerator of a fraction is 5 less than its denominator. If 6 is added to the numerator and 4 to the denominator, the fraction is doubled. What is the fraction?

Answer by mananth(16946)

The numerator of a fraction is 5 less than its denominator.
Let numerator be x
denominator will be x+5
The fraction is x/(x+5)

If 6 is added to the numerator and 4 to the denominator, the fraction is doubled.
(x+6)/(x+5+4)= 2* (x/(x+5))
(x+6)/(x+9) = 2x/(x+5)
(x+6)(x+5) =2x(x+9)
x^2+11x+30 =2x^2+18x
Re arrange
x^2+7x-30 =0
x^2+10x-3x-30=0
x(x+10)-3(x+10)=0
(x+10)(x-3)=0
x=-10 or 3
Now write the fraction